A tree with a height of 12 yards casts a shadow that is 33 yards long at a certain time of day. At the same time, another tree nearby casts a shadow that is 20 yards long. How tall is the second tree?
Answer by Cromlix(4381)

Similar triangles.
Height of first tree/ x (Unknown height of other tree)
= Length of first shadow/ Length of second shadow
12/x = 33/20
Cross multiply
33x = 240
x = 240/33
x = 7.27 yds.
This is the height of the second tree.
Hope this helps.
:-)